explain why F = 9 over 5C + 32
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(57384)

explain why F = 9 over 5C + 32
------------
That equation relates Farenheit and Celcius termperature scoles.
----
Freezing: C = 0 and F = 32
Boiling: C - 100 and F = 212
-----------------------------------
Using those points you get:
slope = (212-32)/(100-0) = 180/100 = 9/5
----------------
Since you have the point (0,32) you have the intercept
-----
So the equation is F = (9/5)C + 32
==============================================
Cheers,
Stan H.

Answer by Alan3354(30993)

http://david.swaim.com/tempconv.htm